OpenStudy (abbycross167):

Can someone please help me answer this math question? Please? I'll fan and give a medal!! I'm working on polynomial division Question - Antonio is playing a watermelon seed spitting contest with his siblings. Not only do they want to know who can spit the seed the farthest, but also whose seed stays in the air the longest. The formula for Antonio’s seed height h in feet after t seconds is: h=-16t2 + 8t + 6. Since the height of the seed is 0 when it hits the ground, the equation becomes -16t2 + 8t + 6 = 0. Make the equation simpler to solve by dividing both sides by the common factor. Solution: -16t^2 + 8t + 6 = 0 Common factor = 6 -16t^2 + 8t + 6 = 0 ---------------------- 6 6 -16t^2 + 8t + 6 = 0 ------ ----- ---- 6 6 6 -0.3752 + .75 + 1 = 0

OpenStudy (sweetnikkie072):

question?

OpenStudy (abbycross167):

@sweetnikkie072 here you go... sorry about that

OpenStudy (sweetnikkie072):

do you know factorizing using middle term splitting

OpenStudy (abbycross167):

umm no ma'am/sir I dont think I have ever heard of that

OpenStudy (sweetnikkie072):

okhay..so is it important to be in decimals coz if u want leave it in whole numbers...let the common factor be 2

OpenStudy (sweetnikkie072):

-8t^2+4t+3

OpenStudy (sweetnikkie072):

does that work?

jhonyy9 (jhonyy9):

-16t^2 + 8t + 6 = 0 /6 -2(8t^2 -4t -3) -------------- = 0 6 -(8t^2 -4t -3) -------------- =0 3